OpenStudy (anonymous):

write an equation in slope-inyercept form for the line parallel to y=5x-2 that passes through the point (8,-2)

OpenStudy (anonymous):

OK, so first you are going to figure out the slope of your new line, which being parallel to line y = 5x -2 would give your line a slope of 5. Next you put your equation into point slope form\[y - y _{1} = m (x -x _{1})\] y+2 = 5(x-8) and simplify y + 2 = 5x -40 Then you convert it to slope intercept y = mx + b y = 5x -42 This should be your answer

OpenStudy (anonymous):

Since we have the slope, which is 5 just use the y intercept form and plugin your points y = mx + b -2 = 5(8) + b -2 = 40 +b -40-2 = b -42 = b Now that we have b, which is the y intercept point we can write the equation of the line in slope intercept form y = 5x - 42

OpenStudy (anonymous):

good ^